FOURIER TRANSFORM MICROWAVE SPECTRUM OF THE AlC_{2} (X̃^{2}A_{1}) RADICAL

\underline{D. T. HALFEN}, J. MIN, and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of the AlC_{2} (X̃^{2}A_{1}) radical in the range 4 - 60 GHz has been measured using Fourier transform microwave (FTMW) methods. The species was produced using the Discharge Assisted Laser Ablation Source (DALAS) technique in a supersonic jet expansion of aluminum vapor and CH_{4}, diluted in argon carrier gas. The N = 1 → 0 and 2 → 1 transitions have been measured near 22 and 44 GHz, each exhibiting fine structure and hyperfine splittings, arising from the nuclear spin of aluminum, I(^{27}Al) = 5/2. The higher frequency transition was recorded using a newly constructed U-band FTMW system operating at 40 - 60 GHz. The data have been analyzed with a case (b) asymmetric top Hamiltonian, and rotational, fine structure, and hyperfine constants have been determined. Measurements of the ^{13}C isotopologues are currently underway to establish a precise structure for AlC_{2}.